Misacc:

Gabriel used the expression 2.5x + 2y – 2 to represent his total cost, and 28 – 2.5x – 2y to represent the amount of change he should receive from $30. What was Gabriel’s error?

jimthompson5910:

A = total cost = 2.5x+2y - 2 B = amount of money handed over = 30 C = amount in change he gets back C = B - A C = 30 - (2.5x+2y - 2) C = ??

Misacc:

Thank you

jimthompson5910:

Let me know what you get when you simplify 30 - (2.5x+2y - 2)

Misacc:

Im stupid soooooo idk but I got an answer you just needded to add the -2 to 30

jimthompson5910:

Let's simplify the expression 30 - (2.5x+2y - 2) 30 - 2.5x-2y + 2 .... distribute the negative through to each term inside 30+2-2.5x-2y 32-2.5x-2y So the mistake Gabriel made was subtracting 2, instead of adding 2. I think that's what you meant to say?

Misacc:

The entire expression should be subtracted from 30. To subtract the quantity, distribute the negative sign, which changes the signs on the terms. So, the –2 changes to positive and gets added to 30 instead of subtracting from it. was the sample answer

jimthompson5910:

Yes that answer works.

pretty123:

Do yall know this A store purchases an item wholesale for $60 and marks it up by 70 percent. If there is a 7 percent tax on the item, how much will the item cost?

jimthompson5910:

If the store marks it up by 70%, then you multiply by 1.70 to get 60*1.70 = 102 So the item was $60 but now it's $102 Now we add on the 7% tax So we multiply by 1.07 to do this 1.07*102 = 109.14 The final cost after tax is $109.14

pretty123:

thanks

jimthompson5910:

No problem